# Sparse but Wrong: Incorrect L0 Leads to Incorrect Features in Sparse Autoencoders

### v2

@@ A.3 LLM SAE training details
For JumpReLU SAEs, we broadly follow the training procedure laid out by Conerly et al. (2025). However, we do not apply learning rate decay, and only warm $\lambda_{s}$ for 100M tokens to avoid the sparsity penalty changing throughout the majority of training. We use a learning rate of $2e^{-4}$, $c=4$, $\lambda_{p}=3e-6$ and bandwidth $\epsilon=2.0$ as recommended by (Conerly et al., 2025).
A.4 Transitioning L0 during training
@@
A.5 Alternative metric: decoder pairwise cosine similarity
We find that $n$th decoder projection is not the only metric that works to gauge the extent to which features are being mixed across SAE latents. Simply taking the mean cosine similarity of each decoder latent with every other decoder latent seems to work as well. We call this metric decoder pairwise cosine similarity, $c$, and define it as below:
$c=\frac{1}{\binom{h}{2}}\sum_{i=1}^{h-1}\sum_{j=i+1}^{h}|\cos(\mathbf{W}_{\text{dec},i},\mathbf{W}_{\text{dec},j})|$ (6)
where $\binom{h}{2}=\frac{h(h-1)}{2}$ is the total number of distinct pairs.
The intuition behind this metric is that if SAE decoder latents are mixing lots of positive and negative components of correlated and anti-correlated features, then each SAE latent should become less orthogonal to each other SAE latent, as many latents will likely mix together similar features. This should mean that the absolute value of the cosine similarity between arbitrary latents should also increase the worse this mixing becomes.
A.5.1 Toy model results
We calculate pairwise calculate similarity $c$ for each of the BatchTopK SAEs we trained on toy models from Section 3.5. Results are shown in Figure 13. We see that pairwise cosine similarity is minimized at the true L0.
Figure 13: Decoder pairwise cosine similarity evaluated on 5 seeds of toy model SAEs. The true L0 is indicated with a dotted line at 11. The shaded area is 1 stdev. The metric is minimized at the true L0.
A.5.2 LLM SAE results
@@
Next, we calculate pairwise cosine similarity for each LLM SAE we evaluated in the paper along with k=16 sparse probing results. Gemma-2-2b layer 5 and Llama-3.2-1b layer 7 results are shown in Figure 14. The results roughly match what we saw with $s_{n}^{\text{dec}}$, being minimized at perhaps a slightly lower but still similar L0.
@@
Figure 14: Decoder pairwise cosine similarity metric and k=16 sparse probing results for BatchTopK SAEs trained on Gemma-2-2b layer 5 (left) and Llama-3.2-1b layer 7 (right). The metric is roughly minimized near peak sparse-probing performance. The shaded area is 1 stdev.
@@
Figure 15: Decoder pairwise cosine similarity metric and sparse probing results for BatchTopK and JumpReLU SAEs trained on Gemma-2-2b layer 12. The metric seems to align less well with k=16 sparse probing results (left), but aligns reasonably well with k=1 sparse probing results (right). The pairwise cosine similarity plot is repeated on both sides to make a vertical comparison with the k-sparse probing results easier.
BatchTopK and JumpReLU results for Gemma-2-2b layer 12 are shown in Figure 15. Interestingly, the BatchTopK results seem to be minimized at a slightly earlier L0 than the JumpReLU results under this metric, which is different than the $n$th decoder projection results from Section 4, where we saw both minimized at roughly the same point. We note that the decoder pairwise cosine similarity metric seems to align better with k=1 sparse probing results, while $n$th decoder projection seems to align better with k=16 sparse probing results. We do not know if this is a meaningful difference between these metrics. It is possible that, as we saw in Section 4.2, there is a range of L0 values that can be too high for some latents while too low for other latents in BatchTopK SAEs, and perhaps the difference between metrics (and k=1 vs k=16 sparse probing results) relates to differences in sensitivity to this mixture of too high / too low latents. Regardless, in all cases, when L0 is low both metrics show large values, indicating high polysemanticity in SAE latents.
A.5.3 Which metric is better?
We choose to focus on $n$th decoder projection in this paper since we feel it allows for more insightful interpretations of the decoder projection plots (e.g. Section 4.2). However, we do not feel that either metric is inherently better or worse than the other, and we do not have a strong reason to pick either metric. We expect that when an SAE is near the correct L0, there are likely many indicators that all should point to similar results. We thus recommend that practitioners try both metrics to see which results in easier interpretation for their use-case.
A.6 Automatically finding the correct L0 during training
@@
We then provide this gradient to the Adam optimizer (Kingma & Ba, 2014) with default settings, and allow it to change the L0 parameter.
@@
A.7 L0 of open-source SAEs on Neuronpedia
We analyze common open-source SAEs as provided by Neuronpedia (Lin, 2023) and SAELens (Bloom et al., 2024). We include all SAEs cross-listed in both SAELens and Neuronpedia with an L0 reported in SAELens. We show the results as a histogram in Figure 16. Our analysis shows that for layer 12 of Gemma-2-2b, the correct L0 should be around 200-250. However, we find that most open-source SAEs have L0 below 100, much lower than our analysis expects to be ideal.
@@ A.7 L0 of open-source SAEs on Neuronpedia
A.8 Limitations
We limited the scope of our investigation to features satisfying the linear representation hypothesis, and do not investigate how SAEs react if the underlying features are actually non-linear (Engels et al., 2025). However, we do not feel that non-linear features are necessary for SAEs to fail to work properly, as we demonstrate in this paper. We also do not consider the nuances of how unbalanced correlations impact the SAE, as simple correlations are already enough to cause problems. However, we do expect that different sorts of correlations may affect SAEs differently, and would encourage future work to look into this. Finally, we only investigated a few layers of popular LLMs, as running sweeps of SAE training at every layer of the LLM was too prohibitively expensive for this work. Nevertheless, we have no reason to expect any meaningfully different behavior in decoder projection at other LLM layers.
A.9 Extended Nth decoder projection plots
@@
A.10 Extended analysis of JumpReLU vs BatchTopK dynamics
JumpReLU and BatchTopK SAEs are both considered state of the art, but we find they have notable differences in their behavior at high L0 in our experiments. In this section, we explore what maybe be causing these differences. In theory, JumpReLU and BatchTopK SAEs are very similar, as a BatchTopK SAE can be viewed as a JumpReLU SAE with a single global threshold, rather than a threshold per-latent (Bussmann et al., 2024). However, the training losses are quite different for JumpReLU vs BatchTopK. We use the JumpReLU variant laid out by Conerly et al. (2025), which allows gradients to flow through the JumpReLU threshold to the rest of the model parameters. We expect this means that JumpReLU SAEs are better able to coordinate the threshold with the rest of the model parameters, while BatchTopK cannot, as the threshold does not directly receive a gradient in BatchTopK training.

### v3

@@ A.7 LLM SAE training details
For JumpReLU SAEs, we broadly follow the training procedure laid out by conerly2025jumprelu. However, we do not apply learning rate decay, and only warm $\lambda_{s}$ for 100M tokens to avoid the sparsity penalty changing throughout the majority of training. We use a learning rate of $2e^{-4}$, $c=4$, $\lambda_{p}=3e^{-6}$ and bandwidth $\epsilon=2.0$ as recommended by (conerly2025jumprelu).
A.8 Transitioning L0 during training
@@
A.9 Alternative metric: nth decoder projection
Figure 17: Idealized histogram of decoder projections on input activations demonstrating the intuition behind our $n$th decoder projection metric, $s_{n}^{\text{dec}}$. For an arbitrary input, most latents should be non-active and thus have low projection. When SAE latents are monosemantic, meaning they do not mix components of many features, we expect non-active latents to have a near-zero projection on arbitrary input activations. However, if SAE latents mix positive and negative components of many underlying features, then those latents will have larger projections on arbitrary inputs that contains those features. By picking an NN less than $h/2$ (corresponding roughly to the origin), a smaller $s_{n}^{\text{dec}}$ means latents have smaller projection on arbitrary inputs and thus are more monosemantic.
$s_{n}^{\text{\text{dec}}}=\mathbf{z}_{\downarrow}[n\cdot b]$ (59)
where $n\cdot b]$ corresponds to selecting the element at index $n\cdot b$. The multiplication by $b$ accounts for the batch dimension, effectively selecting the $n$th highest projection value when considering all samples in the batch. For this to work $n$ should be sufficiently larger than a reasonable guess at the correct L0, as in a perfect SAE, the decoder for these latents should be uncorrelated with input activations. Picking any $n$ up to $\nicefrac{{h}}{{2}}$ should work, as the majority of latents should have low projection on arbitrary input activations, so $\nicefrac{{h}}{{2}}$ intuitively corresponds to 0 expected projection. The intuition behind $s_{n}^{dec}$ is shown visually in Figure 17, and is formalized in Appendix A.10.
A.9.1 LLM SAE results
@@
Next, we $s_{n}^{\text{dec}}$ for each LLM SAE we evaluated in the paper along with k=16 sparse probing results. Gemma-2-2b layer 5 and Llama-3.2-1b layer 7 results are shown in Figure 19. The results roughly match what we saw with $c_{\text{dec}}$.
@@
Figure 19: $n$th decoder projection and k=16 sparse probing results for BatchTopK SAEs trained on Gemma-2-2b layer 5 (left) and Llama-3.2-1b layer 7 (right). The metric is roughly minimized near peak sparse-probing performance. The shaded area is 1 stdev.
@@
Figure 20: $n$th decoder projection and sparse probing results for BatchTopK and JumpReLU SAEs trained on Gemma-2-2b layer 12. The metric seems to align less well with k=16 sparse probing results.
BatchTopK and JumpReLU results for Gemma-2-2b layer 12 are shown in Figure 20. The results look similar to what we saw for $c_{\text{dec}}$.
A.9.2 Which metric is better?
We choose to focus on $c_{\text{dec}}$ as it is a simpler metric both to understand and implement as it has no hyperparameters. However, we expect that when an SAE is near the correct L0, there are likely many indicators that all should point to similar results. Any metric which can detect correlated features being mixed into SAE latents should give roughly similar results.
A.10 Theoretical Justification for $s_{n}^{\text{dec}}$ Metric
We provide a theoretical justification for why the $n$th decoder projection metric, $s_{n}^{\text{dec}}$, successfully identifies when SAE latents are mixing correlated features.
Theorem 3.
Consider two SAEs with identical dictionary sizes $h$ and input dimension $d$, where SAE 1 has greater feature mixing than SAE 2. Specifically, let the decoder projections onto a feature $f$ for non-active latents follow:
SAE 1: $\quad z_{i}^{(1)}\sim\mathcal{N}(0,\sigma_{0}^{2}+\sigma_{1}^{2})$ (60)
SAE 2: $\quad z_{i}^{(2)}\sim\mathcal{N}(0,\sigma_{0}^{2}+\sigma_{2}^{2})$ (61)
where $\sigma_{0}^{2}$ represents the base variance from superposition noise, and $\sigma_{1}^{2}>\sigma_{2}^{2}$ represents the variance from feature mixing. Then for $n<h/2$, we have $\mathbb{E}[s_{n}^{\text{dec}}]$ is larger for SAE 1 than SAE 2.
Proof.
Let $f\in\mathbb{R}^{d}$ be an underlying true feature with $\|f\|_{2}=1$. Consider an SAE with decoder $\mathbf{W}_{\text{dec}}\in\mathbb{R}^{d\times h}$ and decoder bias $\mathbf{b}_{\text{dec}}\in\mathbb{R}^{d}$. For an input activation $\mathbf{x}\in\mathbb{R}^{d}$, the projection of latent $i$ onto the input is:
$z_{i}=(\mathbf{x}-\mathbf{b}_{\text{dec}})^{\top}\mathbf{W}_{\text{dec},i}$ (62)
Decomposition of decoder latents.
We decompose each decoder latent $\mathbf{W}_{\text{dec},i}$ into three components:
$\mathbf{W}_{\text{dec},i}=\alpha_{i}f+\beta_{i}g_{i}+\epsilon_{i}$ (63)
where:
  • $\alpha_{i}f$ is the component aligned with feature $f$ (the intended feature for latent $i$ if $i$ is the correct latent, or mixing if $i$ is incorrect)
  • $\beta_{i}g_{i}$ represents components of other correlated/anti-correlated features mixed into latent $i$, where $g_{i}$ is orthogonal to $f$
  • $\epsilon_{i}$ represents superposition noise, also orthogonal to $f$
Distribution of projections for non-active latents.
Consider latents that should not activate for feature $f$ (i.e., latents $i$ where $\alpha_{i}$ should ideally be near zero). For an input $\mathbf{x}$ containing feature $f$ with magnitude $m_{f}$, we can write:
$\mathbf{x}-\mathbf{b}_{\text{dec}}=m_{f}f+\mathbf{r}$ (64)
where $\mathbf{r}$ contains all other feature contributions orthogonal to $f$.
The projection of latent $i$ becomes:
$z_{i}=(m_{f}f+\mathbf{r})^{\top}(\alpha_{i}f+\beta_{i}g_{i}+\epsilon_{i})=m_{f}\alpha_{i}+\mathbf{r}^{\top}(\beta_{i}g_{i}+\epsilon_{i})$ (65)
For non-active latents in a well-trained SAE, we expect:
  • $\alpha_{i}\approx 0$ for the intended feature component
  • $\beta_{i}g_{i}$ represents unintended mixing of correlated features
  • $\epsilon_{i}$ represents superposition noise
Modeling as Gaussian mixtures.
Under the assumptions that:
  1. 1.
    Feature magnitudes $m_{f}$ and residual components $\mathbf{r}$ vary across the input distribution
  2. 2.
    The number of latents $h$ is large
  3. 3.
    Feature mixing coefficients $\beta_{i}$ arise from optimization pressure to compensate for insufficient L0
By the Central Limit Theorem, the distribution of projections $z_{i}$ for non-active latents approximately follows:
$z_{i}\sim\mathcal{N}(0,\sigma_{\text{base}}^{2}+\sigma_{\text{mix}}^{2})$ (66)
where:
  • $\sigma_{\text{base}}^{2}$ captures variance from superposition noise ($\epsilon_{i}$)
  • $\sigma_{\text{mix}}^{2}$ captures variance from feature mixing ($\beta_{i}g_{i}$)
Comparing two SAEs.
Consider two SAEs:
  • SAE 1 (high feature mixing): $z_{i}^{(1)}\sim\mathcal{N}(0,\sigma_{\text{base}}^{2}+\sigma_{1}^{2})$ where $\sigma_{1}^{2}$ is large
  • SAE 2 (low feature mixing): $z_{i}^{(2)}\sim\mathcal{N}(0,\sigma_{\text{base}}^{2}+\sigma_{2}^{2})$ where $\sigma_{2}^{2}$ is small
with $\sigma_{1}^{2}>\sigma_{2}^{2}$.
Computing $s_{n}^{\text{dec}}$.
For a batch of size $b$, we have $bh$ projection values. After sorting in descending order, the $n$th decoder projection is:
$s_{n}^{\text{dec}}=z_{\downarrow}[n\cdot b]$ (67)
This corresponds to the $(n\cdot b)/(bh)=n/h$ quantile of the distribution. For $n<h/2$, this is the $(n/h)$th quantile on the positive side of the distribution.
Quantile comparison.
For a standard normal distribution $Z\sim\mathcal{N}(0,1)$ and $\sigma_{1}>\sigma_{2}>0$, the $p$th quantile satisfies:
$Q_{p}(\mathcal{N}(0,\sigma_{1}^{2}))=\sigma_{1}\cdot Q_{p}(\mathcal{N}(0,1))>\sigma_{2}\cdot Q_{p}(\mathcal{N}(0,1))=Q_{p}(\mathcal{N}(0,\sigma_{2}^{2}))$ (68)
for $p>0.5$ (corresponding to positive quantiles).
Since $n<h/2$ implies $p=n/h<0.5$, we are actually looking at the (1p)(1-p)th quantile on the right tail due to descending sort. This gives us:
$\mathbb{E}[s_{n}^{\text{dec}}]_{\text{SAE 1}}>\mathbb{E}[s_{n}^{\text{dec}}]_{\text{SAE 2}}$ (69)
Therefore, SAEs with greater feature mixing (larger $\sigma_{\text{mix}}^{2}$) will have larger values of $s_{n}^{\text{dec}}$, justifying its use as a metric for detecting feature mixing. ∎
Remark 2.
The choice of $n\approx h/2$ in practice corresponds to sampling from a region where the distribution is sensitive to changes in variance (roughly near the median), while being sufficiently far from the extreme tails to maintain statistical stability. Values of $n$ too close to 0 would sample from the extreme right tail where variance is high, while $n$ too close to $h$ would sample from regions dominated by active latents rather than the non-active latents we wish to characterize.
Remark 3.
This theoretical analysis assumes that decoder projections follow approximately Gaussian distributions. While this is a simplification, our empirical results in both toy models (where we have full control) and LLM SAEs support this assumption, as evidenced by the decoder projection histograms in Figure 9.
A.11 Automatically finding the correct L0 during training
@@
We then provide this gradient to the Adam optimizer (kingma2014adam) with default settings, and allow it to change the L0 parameter.
@@
A.12 Extended LLM results
We include further results for Gemma-2-2b layer 20, to extend the analysis to later model layers. Results are shown in Figure 21.
Figure 21: Nth decoder projection (top left) and decoder pairwise cosine similarity (top right) with K=16 sparse probing results (bottom-left) and K=1 sparse probing results (bottom-right) for BatchTopK SAEs trained on Gemma-2-2b layer 20.
A.13 L0 of open-source SAEs on Neuronpedia
We analyze common open-source SAEs as provided by Neuronpedia (neuronpedia) and SAELens (bloom2024saetrainingcodebase). We include all SAEs cross-listed in both SAELens and Neuronpedia with an L0 reported in SAELens. We show the results as a histogram in Figure 22. Our analysis shows that for layer 12 of Gemma-2-2b, the correct L0 should be around 200-250. However, we find that most open-source SAEs have L0 below 100, much lower than our analysis expects to be ideal.
@@ A.13 L0 of open-source SAEs on Neuronpedia
A.14 Limitations
We limited the scope of our investigation to features satisfying the linear representation hypothesis, and do not investigate how SAEs react if the underlying features are actually non-linear (engels2025not). However, we do not feel that non-linear features are necessary for SAEs to fail to work properly, as we demonstrate in this paper. We also do not consider the nuances of how unbalanced correlations impact the SAE, as simple correlations are already enough to cause problems. However, we do expect that different sorts of correlations may affect SAEs differently, and would encourage future work to look into this. Finally, we only investigated a few layers of popular LLMs, as running sweeps of SAE training at every layer of the LLM was too prohibitively expensive for this work. Nevertheless, we have no reason to expect any meaningfully different behavior in decoder projection at other LLM layers.
A.15 Extended Nth decoder projection plots
@@
A.16 Extended analysis of JumpReLU vs BatchTopK dynamics
JumpReLU and BatchTopK SAEs are both considered state of the art, but we find they have notable differences in their behavior at high L0 in our experiments. In this section, we explore what maybe be causing these differences. In theory, JumpReLU and BatchTopK SAEs are very similar, as a BatchTopK SAE can be viewed as a JumpReLU SAE with a single global threshold, rather than a threshold per-latent (bussmann2024batchtopk). However, the training losses are quite different for JumpReLU vs BatchTopK. We use the JumpReLU variant laid out by conerly2025jumprelu, which allows gradients to flow through the JumpReLU threshold to the rest of the model parameters. We expect this means that JumpReLU SAEs are better able to coordinate the threshold with the rest of the model parameters, while BatchTopK cannot, as the threshold does not directly receive a gradient in BatchTopK training.
@@ A.16 Extended analysis of JumpReLU vs BatchTopK dynamics
A.17 Pytorch pseudocode for metrics
def pairwise_decoder_cosine_similarity(sae):
    norm_dec = torch.nn.functional.normalize(sae.W_dec, dim=1)
    dec_sims = torch.mm(norm_dec, norm_dec.T)
    triu_mask = torch.triu(
        torch.ones_like(dec_sims),
        diagonal=1,
    ).bool()
    return dec_sims[triu_mask].abs().mean()
Figure 28: Pytorch pseudocode for decoder pairwise cosine similarity
def nth_decoder_projection(input_acts, sae, n):
    dec_proj = (input_acts - sae.b_dec) @ sae.W_dec.T
    sorted_dec_proj = dec_proj.flatten().sort(descending=True)
    index = n * dec_proj.shape[0]
    return sorted_dec_proj.values[index]
Figure 29: Pytorch pseudocode for nth decoder projection
We present Pytorch pseudocode for nth decoder projection in Figure 29 and decoder pairwise cosine similarity in Figure 28.